Now, assume that Hillary’s annual income changes from $55,000 to $60,000 and this

promtps a change in the number of 16oz pumpkin lattes she consumes annually. Find the

number of lattes that Hillary consumes after the income change, if she used to consume

360 lattes before experiencing an income change, and the income elasticity is 1.25.
Answer #1

% change in income = (60,000 - 55,000)/55,000 * 100 = 5,000/55,000 * 100 = 1/11*100 = 9.09%

Income elasticity of demand = % change in quantity / % change in income

1.25 = % change in quantity/9.09%

% change in quantity = 1.25 * 9.09 = 11.3625 = 11 lattes

Therefore, after the income increase, total consumption of latte would be = 360 + 11 = 371.